Lena Tucker contributed $20,000 in cash and land with an adjusted basis of $100,000 and a fair market value of $140,000 to a newly formed corporation in return for 1,000 shares of stock in the corporation. During Year 1, Lena’s pro rata share of the corporation's separately and nonseparately stated items of income was $40,000 (including $5,000 of tax-exempt interest), and her pro rata share of the corporation's separately stated items of loss and deductions was $80,000. During the year she received distributions totaling $30,000. What is her adjusted basis in the 1,000 shares of stock?
Answer #1

Adjusted basis is a concept driven by the IRS code, which requires a tax payer to calculate the adjusted cost of assets for the purposes of tax accounting and declaration of accurate income for the purposes of tax payments.

In general terms, adjusted cost can be understood as a mechanism to calculate the as on date value of assets considering initial investment or purchase cost (including all other ancillary cost such as brokerage, commission, etc), investment post acquisition (For example re-investment of dividend), depreciation or amortization and any other monitory event increasing or decreasing the value of assets.

For calculating adjusted cost, following formula is considered:

1. Purchase Cost

2. Add: Re-investments

3. Less: Depreciation or amortization

4. Adjusted cost = 1+2-3

Considering the above concept, adjusted cost of shares for Lena shall be as below:

1. Purchase cost of shares = Cash payment of $20,000 + Adjusted cost of land of $100,000 which is equal to $120,000

2. Appreciation = Lena’s pro rata share of the corporation's separately and nonseparately stated items of income was $40,000

3. Depreciation = Pro rata share of the corporation's separately stated items of loss and deductions was $80,000 + Distributions totaling $30,000 = $130,000

Adjusted Cost of stock = 1 + 2 - 3 = $120,000 + $40,000 - $130,000 = $30,000